OpenStudy (anonymous):

Veronica teaches French in groups of four students. She rents a room for $200 per month. She charges $45 per hour per student. Veronica teaches seven groups of students, and her net earnings per month are $6,100. Part A: If each student learns French for the same number of hours, write an equation that can be used to calculate the number of hours each student learns French in a month. (5 points) Part B: How many hours did each student learn French each month? Show your work and justify each step of your work. (5 points)

OpenStudy (lucaz):

7 groups of 4 students = total of 28 students \[$6100=28\times$45(h)-$200\]

OpenStudy (anonymous):

is the first one the answer to part a

OpenStudy (anonymous):

are you there?

OpenStudy (lucaz):

?

OpenStudy (anonymous):

sorry im just confused

OpenStudy (lucaz):

that was the part A, for part B you solve the equation for h to find the number of hours

OpenStudy (lucaz):

you have a total of $6100, so the equation is (something)=6100

OpenStudy (lucaz):

ew know she pays $200 for the room, then (something)-200=6100

OpenStudy (anonymous):

so i put Part A: 6100=28x45(h)-200 and then can you help me on Part B

OpenStudy (lucaz):

there are 7 groups of 4 students each, 7(4)=28, total of 28 sutdents. each student pays $45 per hour = 45(h) the equation becomes \[28\times45(h)-200=6100\]

OpenStudy (anonymous):

so thats part a?

OpenStudy (lucaz):

yes, thats the equation for the problem, to find the number of hours (h) you solve the equation for h, like you solve any equation for x

OpenStudy (anonymous):

ok i get it thank you dont go anywhere please i am having some trouble with a couple